OpenStudy (anonymous):

can anyone explain this to me? 28(x+5)-70(x+5)^2=[14(x+5)] [2-5(x+5)]

OpenStudy (anonymous):

where do get the 2-5 from?

OpenStudy (anonymous):

To make it a little more readable, let's do this: y=x+5 So: 28(x+5)-70(x+5)^2 ; we start with the left side of the eqn. =28y-70y^2 ; use y=x+5 =(28y-70y^2) =14 * ( 2y - 5y^2 ) ; take 14 out of the parantheses =14y * ( 2 - 5y ) ; take y out of the parantheses =[14(x+5)] [2-5(x+5)] ; and we ended up with the right side of eqn.